# Neurosymbolic Association Rule Mining from Tabular Data

### v1

@@ 4.2 Experimental Setting 2: Aerial+ on Downstream Tasks
The goal of this experimental setting is to evaluate Aerial+ on downstream classification tasks within established rule-based interpretable models.
@@
Table 5 shows the experimental results including the number of rules (CBA) or frequent itemsets (BRL and CORELS), accuracy, and execution times. [^3] The execution time includes rule mining (including training for Aerial+) and classifier construction time. The results show that with a significantly smaller number of rules with Aerial+, rule-based classifiers run substantially faster than with the rules from FP-Growth. Despite having a significantly lower number of rules, all of the rule-based classifiers with rules (or frequent itemsets) from Aerial+ lead to a higher or comparable accuracy on all datasets.
Hyperparameter analysis of Aerial+. We also conducted a hyperparameter analysis for Aerial+ to explore how antecedent and consequent similarity thresholds ($\tau_{a}$ and $\tau_{c}$) affect rule quality. The analysis can be found in Appendix B.
@@ 5 Conclusions
Extensive rule quality evaluations in Section 4.1 on five real-world tabular datasets show that Aerial+ learns a compact set of high-quality association rules with full data coverage while outperforming state-of-the-art methods in execution time on high-dimensional datasets. When used in downstream classification tasks within rule-based interpretable machine learning models, as described in Section 4.2, the concise set of rules learned by Aerial+ significantly reduces execution time while maintaining or improving accuracy. Aerial+ supports parallel and GPU execution (Section 3.3) and both evaluations on rule quality and downstream tasks, together with the runtime complexity analysis, indicate that Aerial+ is scalable on large datasets. In addition, two variants of Aerial+ can perform ARM with item constraints and frequent itemset mining. We also further outline how other solutions to rule explosion can be integrated into Aerial+.
@@
Acknowledgement. This work has received support from The Dutch Research Council (NWO), in the scope of the Digital Twin for Evolutionary Changes in water networks (DiTEC) project, file number 19454.
@@ Appendix A Runtime Complexity Analysis of Aerial+
Lines 2-14 iterates over the number of antecedents $a$ for the outer loop, meaning the outer loop runs $O(a)$ times.
@@ Appendix C Variations of Aerial+
This section presents two variations to Aerial+’s rule extraction method given in Algorithm 1, and further describes how other ARM variations such as top-k rule mining [Fournier-Viger et al., 2012] can be incorporated into Aerial+. Modifications in the proposed variations relative to Algorithm 1 are distinguished using a light yellow background.

### v2

@@ 4.2 Experimental Setting 2: Aerial+ on Downstream Tasks
This setting evaluates Aerial+ on rule-based classification tasks.
@@
Table 5 shows the experimental results including the number of rules (CBA) or frequent itemsets (BRL and CORELS), accuracy, and execution times. [^3] The execution time includes rule mining (including training for Aerial+) and classifier construction time. The results show that with a significantly smaller number of rules (or frequent itemsets) with Aerial+, rule-based classifiers run substantially faster than with the rules from FP-Growth. Despite having a significantly lower number of rules, all of the rule-based classifiers with rules (or frequent itemsets) from Aerial+ lead to a higher or comparable accuracy on all datasets.
Hyperparameter analysis. Appendix B (omitted here for brevity) shows how similarity thresholds ($\tau_{a}$, $\tau_{c}$) affect Aerial+ rule quality.
Qualitative advantages of Aerial+. Aerial+ offers two qualitative advantages over the state-of-the-art: i) once trained in O(n) time, it can verify whether a given association exists by creating the corresponding test vector in O(1) time—unlike other methods, ii) it can be integrated into larger neural networks for interpretability.
@@ 5 Conclusions
Extensive rule quality evaluations in Section 4.1 show that Aerial+ learns a compact set of high-quality association rules with full data coverage, outperforming state-of-the-art methods in execution time on high-dimensional datasets. In downstream classification tasks within rule-based interpretable models (Section 4.2), Aerial+’s concise rules significantly reduce execution time while maintaining or improving accuracy. Aerial+ supports parallel and GPU execution (Section 3.3), and results across rule quality, downstream tasks, and runtime complexity demonstrate its scalability on large datasets. Additionally, two Aerial+ variants support ARM with item constraints and frequent itemset mining. We also describe how other solutions to rule explosion can be integrated into Aerial+.
@@
Acknowledgment. This work has received support from the Dutch Research Council (NWO), in the scope of the Digital Twin for Evolutionary Changes in water networks (DiTEC) project, file number 19454.
@@ Appendix A Runtime Complexity Analysis of Aerial+
Lines 2-14 iterate over the number of antecedents $a$ for the outer loop, meaning the outer loop runs $O(a)$ times.
@@ Appendix C Variations of Aerial+
This section presents two variations to Aerial+’s rule extraction method given in Algorithm 1, and further describes how other ARM variations such as top-k rule mining [Fournier-Viger et al., 2012] can be incorporated into Aerial+. Modifications in the proposed variations relative to Algorithm 1 are distinguished using a light yellow background. Python implementations of the two variants described below (and more) can be found as part of the PyAerial package: https://github.com/DiTEC-project/pyaerial.
@@ Appendix E Effect of Minimum Support Threshold on Classification
Appendix F Formal Justification for Aerial+’s Rule Extraction
This section provides a formal justification as to why Aerial+’s rule extraction (Section 3.3) from its Autoencoder architecture (Section 3.2) works in practice.
F.1 Autoencoder Input
Let $\mathcal{X}\subseteq\{0,1\}^{d}$ denote the space of binary feature vectors (consisting of one-hot encodings of categorical features). Let $x\in\mathcal{X}$ be a data point composed of $k$ categorical features, i.e., columns of a table, where each feature $x_{i}$ takes a value in a finite set $\mathcal{A}_{i}$ of cardinality $C_{i}$.
Each categorical feature is represented by a one-hot encoding over $C_{i}$ positions. Thus, the full input vector $x\in\{0,1\}^{d}$ is a concatenation of such one-hot encodings.
The input is corrupted by adding (or subtracting) uniform noise from the interval $[0,0.5]$ to each entry in $\mathcal{X}$, followed by clipping the result to the range $[0,1]$. The corrupted input is represented with $\tilde{\mathcal{X}}$. Formally, $\tilde{x}\in\tilde{\mathcal{X}}$ is given by:
$\tilde{x}=\mathrm{clip}(x+\epsilon),\quad\epsilon\sim\mathcal{U}([-0.5,0.5]^{d}),\quad\mathrm{clip}(z)=\min(\max(z,0),1).$ (1)
F.2 Autoencoder Architecture
An under-complete denoising autoencoder consists of:
  • Encoder: $f:\tilde{\mathcal{X}}\rightarrow\mathbb{R}^{e}$, with $e<d$ (under-complete).
  • Decoder: $g:\mathbb{R}^{e}\rightarrow\mathbb{R}^{d}$.
  • Categorical projection layer: Let $\mathcal{C}_{i}\subset\{1,\dots,d\}$ be the index subset corresponding to the one-hot encoding of feature $x_{i}$. The decoder produces logits $\mathbf{h}_{\mathcal{C}_{i}}\in\mathbb{R}^{C_{i}}$, and applies softmax:
    $\bar{x}_{\mathcal{C}_{i}}=\text{softmax}(\mathbf{h}_{\mathcal{C}_{i}})\in\left\{\mathbf{p}\in\mathbb{R}^{C_{i}}\;\middle|\;p_{j}\geq 0,\ \sum_{j=1}^{C_{i}}p_{j}=1\right\},$ (2)
    where $\bar{x}\in\bar{\mathcal{X}}$ is the reconstructed form of $\mathcal{X}$.
F.3 Training stage
The model is trained to minimize expected binary cross-entropy loss between the initial input $\mathcal{X}$ and reconstruction $\bar{\mathcal{X}}$:
$\mathcal{L}=\mathbb{E}\left[\sum_{i=1}^{n}\text{BCE}(x_{\mathcal{C}_{i}},\bar{x}_{\mathcal{C}_{i}})\right],$ (3)
Each component $\bar{x}_{\mathcal{C}_{i}}^{(c)}$, where $c\in\{1,\dots,C_{i}\}$, denotes the $c$-th entry of the softmax output vector corresponding to feature $x_{i}$. Since the model is trained to minimize the binary cross-entropy between the true one-hot vector $x_{\mathcal{C}_{i}}$ and the predicted distribution $\bar{x}_{\mathcal{C}_{i}}$, this vector approximates the conditional distribution over categories given the corrupted input:
$\bar{x}_{\mathcal{C}_{i}}^{(c)}\approx\mathbb{P}\left(x_{\mathcal{C}_{i}}^{(c)}=1\mid\tilde{x}\right),$ (4)
This approximation is justified by the denoising objective: during training, the model learns to predict the original categorical value from noisy inputs, making the softmax outputs probabilistic estimates of the true class conditioned on the corrupted input.
F.4 Rule Extraction Stage
To determine whether a specific feature-value assignment implies other feature-value activations, we construct a controlled input probe called a test vector and evaluate the decoder’s output response.
Let $x_{i}\in\mathcal{A}_{i}$ be a categorical feature with domain size $C_{i}$, and let $c\in\{1,\dots,C_{i}\}$ be a particular category. We aim to test whether the assignment $x_{i}=c$ implies other specific feature values.
We define an artificial input vector $A\in[0,1]^{d}$ constructed as follows:
  • For the index subset $\mathcal{C}_{i}\subseteq\{1,\dots,d\}$ corresponding to the one-hot encoding of $x_{i}$, set $A_{\mathcal{C}_{i}}^{(c)}=1$ and $A_{\mathcal{C}_{i}\setminus\{c\}}=0$.
  • For all other categorical feature groups $\mathcal{C}_{j}$ with $j\neq i$, assign uniform probability across the classes: for all $c^{\prime}\in\{1,\dots,C_{j}\}$,
    $A_{\mathcal{C}_{j}}^{(c^{\prime})}=\frac{1}{C_{j}}.$
This creates an input vector where only $x_{i}=c$ is deterministically assigned, while all other feature values have uniform probability. No corruption is applied to the test vector.
Forward pass and output interpretation. Next, we feed the input $A$ through the trained autoencoder to obtain output:
$\bar{x}=\text{g}(\text{f}(A))\in[0,1]^{d},$
where the decoder output, function $g$, also includes the categorical projection layer step. For each feature $j$, $\bar{x}_{\mathcal{C}_{j}}\in\mathbb{R}^{C_{j}}$ is a probability vector over its possible categories, derived from the softmax output of the decoder.
Antecedent similarity confirmation. Confirm that the decoded output preserves the input’s forced condition by checking:
$\bar{x}_{\mathcal{C}_{i}}^{(c)}\geq\mathcal{T}_{a},$
where $\mathcal{T}_{a}\in(0,1)$ is an antecedent similarity threshold. This ensures that the model recognizes $x_{i}=c$ as a highly probable input condition.
Consequent discovery. Search the output $\bar{x}$ for all components $\bar{x}_{\mathcal{C}_{j}}^{(c^{\prime})}$ such that:
[…]
where $\mathcal{T}_{c}\in(0,1)$ is a consequent similarity threshold. The set of all such $(x_{j}=c^{\prime})$ pairs forms the set of predicted consequents:
$B=\left\{(x_{j}=c^{\prime})\;\middle|\;\bar{x}_{\mathcal{C}_{j}}^{(c^{\prime})}\geq\mathcal{T}_{c},\ j\neq i\right\}.$
If both the antecedent similarity and the consequent similarity conditions are satisfied, we conclude with the following rule:
$(x_{i}=c)\Rightarrow\bigwedge_{(x_{j}=c^{\prime})\in B}(x_{j}=c^{\prime}).$ (5)
Note that, the same can be applied for multiple features $x_{i}$ on the antecedent side. Since the decoder output approximates conditional probabilities learned during denoising training,
$\bar{x}_{\mathcal{C}_{j}}^{(c^{\prime})}\approx\mathbb{P}(x_{j}=c^{\prime}\mid x_{i}=c),$ (6)
this probing reveals high-confidence associations in the learned conditional distribution.
F.5 Interpretation of the formal justification
Overall interpretation of the formal justification can be summarized as follows:
  • Under-completeness. The encoder maps the corrupted higher-dimensional input $\tilde{x}$ to a lower-dimensional latent space, encouraging the model to capture prominent co-occurrence (association) patterns.
  • Reconstruction objective. Given the noise process in Equation 1, the model is trained to reconstruct the original input $x$, approximating:
    $\bar{x}_{\mathcal{C}_{i}}^{(c)}\approx\mathbb{P}(x_{i}=c\mid\tilde{x}).$
  • Categorical output. The decoder outputs per-feature softmax distributions over categories. Thus, $\bar{x}_{\mathcal{C}_{i}}^{(c)}\in[0,1]$ quantifies the model’s belief that feature $x_{i}$ takes value $c$, conditioned on the input.
  • Probing with test vectors. Rule extraction relies on probing the decoder with a constructed input $A$ that sets $x_{i}=c$ deterministically and others to uniform noise. The resulting output $\bar{x}$ estimates $\mathbb{P}(x_{j}=c^{\prime}\mid x_{i}=c)$ (Equation 6), enabling the discovery of association rules.
Conclusion. A denoising autoencoder trained on the corrupted categorical data learns a conditional distribution over feature values. The decoder’s categorical projection layer approximates $\mathbb{P}(x_{j}=c^{\prime}\mid x_{i}=c)$, making them suitable for rule extraction via direct probing. This forms the basis of Aerial+’s neurosymbolic association rule mining.